# On a problem of Brown, Erdős and Sós

### 06 Dec 2023 v1

@@ 1 Introduction
Next, we settle Conjecture 1.1 for $k=7$. (Our proof does not work when $r=3$ and $t=2$, but this case has been resolved for all $k$ in [2].)
@@
For general odd $k$ we were not able to prove that the limit $\lim_{n\to\infty}n^{-t}f^{(r)}(n;k(r-t)+t,k)$ exists, even when $r$ is large. Nevertheless, arguments reminiscent of Theorem 1.2 yield an upper bound on $\limsup_{n\to\infty}n^{-t}f^{(r)}(n;k(r-t)+t,k)$, which we suspect might be optimal. As evidence, it is indeed optimal for $k=3$, as shown in [4]. As it is not clear that this bound is tight for any $k,t,r$, we do not try very hard to optimise the dependence of $r$ on $k$ and $t$. In particular, we do not provide an explicitly lower bound on $r$ in terms of $k$ and $t$ in the statement.
Proposition 1.5.
Let $k,t,r$ be positive integers such that $k$ is odd, $t\geq 2$, and $r$ is large with respect to $k$ and $t$. Then
$\limsup\limits_{n\to\infty}n^{-t}f^{(r)}(n;k(r-t)+t,k))\leq\frac{1}{t!}\cdot\frac{2}{2\binom{r}{t}-1}\,.$
@@
Section 2 introduces the relevant notation and collects some preliminary results, including our key proposition (Proposition 2.4) needed for the proofs of Theorems 1.3 and 1.4, which are proved in Section 4 and Section 5, respectively. Section 3 provides the proof of Theorem 1.2, Section 6 provides the proof of Proposition 1.5, and finally Section 7 contains some concluding remarks.
@@ 2.2 Density argument
The approach of Delcourt and Postle [2], while proving Conjecture 1.1 for $r=3$, $t=2$ and any $k\geq 2$, relies on the following reduction: they show that in any sufficiently dense $k$-free $3$-graph, it is possible to find a subgraph with almost the same density which is additionally $\ell^{-}$-free for every $\ell\in[2,k-1]$, i.e. having $\ell^{-}$-configurations is ‘inefficient’ for the extremal $k$-free graph. Here we provide another density-type argument. Before stating the result, we introduce some notation. Given an $r$-graph $\mathcal{F}$, define $J(\mathcal{F})$ to be the $t$-graph with $V(\mathcal{F})$ as vertex set and where a $t$-subset $T\subseteq V(\mathcal{F})$ is an edge of $J(\mathcal{F})$ if and only if there exists an $\ell$-configuration for some $\ell\in\left[\lfloor k/2\rfloor\right]$ whose vertex set contains $T$. Observe that, since every edge is a $1$-configuration, $J(\mathcal{F})$ contains the $t$-shadow of $\mathcal{F}$. Therefore $J(\mathcal{F})$ is a supporting $t$-graph of $\mathcal{F}$ (recall Definition 2.1). Moreover, its non-edge girth is greater than $\lfloor k/2\rfloor$.
@@ 2.3 A useful lemma
We show that we can get a subhypergraph of $\mathcal{F}$ which satisfies (P1), (P2) and (P3) by removing $O(n^{t-1})$ edges, which in turn will imply (P4) as well. Observe that, since $\mathcal{F}$ is $k$-free, $\mathcal{F}$ is also $k^{-}$-free. Moreover we show that $\mathcal{F}$ can made $(k-1)^{-}$-free by removing $O(n^{t-1})$ edges. Let $\mathcal{S}$ be a maximal collection of pairwise edge-disjoint $(k-1)^{-}$-configurations of $\mathcal{F}$. If $|\mathcal{S}|>\binom{n}{t-1}$, then there exists a set $T\subseteq V(\mathcal{F})$ of size $t-1$ which is contained in the $(t-1)$-shadow of two $(k-1)^{-}$-configurations $S_{1}$ and $S_{2}$ in $\mathcal{S}$. Let $e\in S_{2}$ be an edge such that $T\subseteq e$. Then $S_{1}\cup\{e\}$ is a $k$-configuration in $\mathcal{F}$, being a collection of $k$ edges spanning at most $[(k-1)(r-t)+t-1]+(r-|T|)=k(r-t)+t$ vertices, a contradiction to $\mathcal{F}$ being $k$-free. Therefore $|\mathcal{S}|\leq\binom{n}{t-1}$ and, by removing from $\mathcal{F}$ all edges of each $S\in\mathcal{S}$, we obtain a subhypergraph $\mathcal{F}_{0}\subseteq\mathcal{F}$ which is $(k-1)^{-}$-free and satisfies $|\mathcal{F}_{0}|\geq|\mathcal{F}|-O(n^{t-1})$.
@@
Before proving the theorem, we prove an inequality and some claims.
@@
We now prove Theorem 1.2, which asserts that $\lim_{n\to\infty}n^{-t}f^{(r)}(n;k(r-t)+t,k)=\frac{1}{t!}\binom{r}{t}^{-1}$, for $k$ even and $r$ sufficiently large in terms of $t$.
@@
In this section we prove Theorem 1.3, asserting that the limit $\lim_{n\to\infty}n^{-t}f^{(r)}(n;5(r-t)+5)$ exists, for $2\leq t<r$. Our proof uses Proposition 2.4. We recall that $J(\mathcal{F})$ is the $t$-graph on $V(\mathcal{F})$ whose edges are $t$-subsets of $\ell$-configurations in $\mathcal{F}$ with $\ell\leq\lfloor k/2\rfloor$ (this is defined above Proposition 2.4).
@@
Observe that, by definition of $J(S)$, we have that $T\in J(S)$ if and only there exists an edge $e\in S$ with $T\subseteq e$ or there exists a $2$-configuration $S^{\prime}$ in $S$ whose vertex set contains $T$. Since in $\mathcal{G}$ any $2$-configuration and any $3^{-}$-configuration are edge-disjoint, we can rule out the second option. Moreover, a set of size $t$ cannot be in more than one edge of $S$ as, otherwise, $S$ would contain a $2$-configuration, which cannot happen for the same reason. Since $S$ has three edges, we conclude that $|J(S)|=3\binom{r}{t}$. Let $T\in J(S)$. That $T\in J(\mathcal{G})$ is obvious, and we aim to show that $T\not\in J(\mathcal{G}^{\prime})$. For that, note that $T\in J(\mathcal{G}^{\prime})$ if and only if there exists an edge $e\in\mathcal{G}^{\prime}$ with $T\subseteq e$ or there exists a $2$-configuration $S^{\prime}$ of $\mathcal{G}^{\prime}$ whose vertex set contains $T$. The first option cannot happen as, otherwise, $S\cup\{e\}$ would be a $4^{-}$-configuration of $\mathcal{G}$, being a collection of four edges spanning at most $(3r-2t-1)+r-t=4(r-t)+t-1$ vertices, a contradiction to $\mathcal{F}_{1}$ being $4^{-}$-free. Similarly, we can rule out the second option as, otherwise, using that $S$ and $S^{\prime}$ are edge-disjoint, $S\cup S^{\prime}$ would be a $5^{-}$-configuration of $\mathcal{G}$, being a collection of five edges spanning at most $(3r-2t-1)+(2r-t)-t=5(r-t)+t-1$ vertices, a contradiction to $\mathcal{F}_{1}$ being $5^{-}$-free.
@@
Before proving the theorem, we prove two inequalities.
@@
We now prove Theorem 1.4, which asserts that the limit $\lim_{n\to\infty}n^{-t}f^{(r)}(n;7(r-t)+t,7)$ exists for $r>t\geq 2$ and $(r,t)\neq(3,2)$. Our proof uses Proposition 2.4.
@@
6 An upper bound on limsup for $k$ odd
We now prove Proposition 1.5, showing that $\limsup_{n\to\infty}n^{-t}f^{(r)}(n;k(r-t)+t,k)\leq\frac{1}{t!}\cdot\frac{2}{2\binom{r}{t}-1}$ for odd $k$ and large $r$. The proof is similar to that of Theorem 1.2.
Proof of Proposition 1.5.
Let $\mathcal{F}$ be any $k$-free $n$-vertex $r$-graph. A $t$-tight component in $\mathcal{F}$ is a collection of edges that can be ordered as $\{e_{1},\ldots,e_{m}\}$ so that $e_{i+1}$ shares at least $t$ vertices with one of $e_{1},\ldots,e_{i}$, for each $i\in[m-1]$. Then $\mathcal{F}$ is a disjoint union of its tight $t$-components, and every component consists of at most $k-1$ edges (as a component on at least $k$ edges contains a $k$-configuration).
For $i\in[2]$, let $\mathcal{F}_{i}$ be the set of the edges of $\mathcal{F}$ which belong to components of size $i$, and let $\mathcal{F}_{3}:=\mathcal{F}\setminus(\mathcal{F}_{1}\cup\mathcal{F}_{2})$. Then the sets $\mathcal{F}_{1},\mathcal{F}_{2},\mathcal{F}_{3}$ partition $\mathcal{F}$. For $i\in[2]$, define $\mathcal{G}_{i}$ to be the $t$-shadow of $\mathcal{F}_{i}$, and define $\mathcal{G}_{3}$ to be the collection of $t$-sets $T$ such that $T\not\in\mathcal{G}_{1}\cup\mathcal{G}_{2}$ and there is a unique component in $\mathcal{F}_{3}$ that contains a $2$-configuration whose vertex set contains $T$. Observe that the sets $\mathcal{G}_{1},\mathcal{G}_{2},\mathcal{G}_{3}$ are pairwise disjoint, so
$|\mathcal{G}_{1}|+|\mathcal{G}_{2}|+|\mathcal{G}_{3}|\leq\binom{n}{t}.$
Write
$\alpha:=\frac{1}{2}\cdot\left(2\binom{r}{t}-1\right).$
We claim that $|\mathcal{G}_{i}|\geq\alpha|\mathcal{F}_{i}|$ for $i\in[3]$. Assuming this holds, we would get
$|\mathcal{F}|=|\mathcal{F}_{1}|+|\mathcal{F}_{2}|+|\mathcal{F}_{3}|\leq\frac{1}{\alpha}\left(|\mathcal{G}_{1}|+|\mathcal{G}_{2}|+|\mathcal{G}_{3}|\right)\leq\frac{2}{2\binom{r}{t}-1}\binom{n}{t}\,.$
from which the result follows.
Notice that $|\mathcal{G}_{1}|=\binom{r}{t}|\mathcal{F}_{1}|\geq\alpha|\mathcal{F}_{1}|$ and $|\mathcal{G}_{2}|=\left(2\binom{r}{t}-1\right)\frac{|\mathcal{F}_{2}|}{2}=\alpha|\mathcal{F}_{2}|$. It thus remains to prove $|\mathcal{G}_{3}|\geq\alpha|\mathcal{F}_{3}|$. For each component $C$ in $\mathcal{F}_{3}$, fix a $2$-configuration of $C$, denoted $S_{C}$, and let $S_{C}^{\prime}$ be a $3$-configuration in $C$ containing $S_{C}$ (these exist since $C$ is a component of size at least three). We now prove the following assertions.
  1. (i)
    The number of $t$-subsets of $V(S_{C})$ which are in the shadow $\partial_{t}(\mathcal{F}\setminus C)$ is at most $(k-3)\binom{2t-2}{t}$;
  2. (ii)
    the number of $t$-subsets of $V(S_{C})$ that are contained in the vertex set of a $2$-configuration $S$ with $S\not\subseteq C$ is at most $(k-2)(k-5)\cdot\binom{4t-4}{t}$.
First observe that, if $e$ and $f$ are edges in different components, then $|e\cap f|<t$ as otherwise $e$ and $f$ would be in the same $t$-component. We will make use of this property several times, without further mentioning it.
Property (i) can be shown as follows. There are at most $k-3$ edges in $\mathcal{F}\setminus C$ that intersect $V(S_{C})$ in at least $t$ vertices (as otherwise, together with $S_{C}$, they would form a $k$-configuration). Each of them intersects $V(S_{C})$ in at most $2(t-1)$ vertices. Therefore, the number of $t$-subsets of $V(S_{C})$ which are in $\partial_{t}(\mathcal{F}\setminus C)$ is at most $(k-3)\cdot\binom{2t-2}{t}$.
For (ii), first notice that if $S$ is a $2$-configuration in a component which is not $C$, then $|V(S)\cap V(S_{C})|\leq 4(t-1)$. Let $\mathcal{S}$ be a maximal collection of pairwise edge-disjoint $2$-configurations that are not contained in $C$ and whose vertex sets intersect $V(S_{C})$ in at least $t$ vertices. Then $|\mathcal{S}|\leq(k-5)/2$, because otherwise the union of $(k-5)/2+1=(k-3)/2$ many $2$-configurations in $\mathcal{S}$ with the $3$-configuration $S_{C}^{\prime}$ forms a $k$-configuration. Now, the number of $2$-configurations $S$ satisfying $S\not\subseteq C$ and $|V(S)\cap V(S_{C})|\geq t$ is at most $(k-2)(k-5)$, because every such $2$-configuration contains an edge $e$ from $\mathcal{S}$, so its other edge is in the same component as $e$, leaving at most $k-2$ options for this other edge. Thus, the number of $t$-subsets of $V(S_{C})$ which are contained in the vertex set of a $2$-configuration $S$ with $S\not\subseteq C$ is at most
$(k-2)(k-5)\cdot\binom{4t-4}{t}\,,$
as claimed in (ii).
It follows from (i) and (ii) that the number of $t$-subsets of $V(S_{C})$ that are in $\mathcal{G}_{3}$ is at least
$\binom{2r-t}{t}-(k-3)\binom{2t-2}{t}-(k-2)(k-5)\cdot\binom{4t-4}{t}\geq k\cdot\left(2\binom{r}{t}-1\right)\geq\alpha|C|\,,$
using that $r$ is large with respect to $t$ and $k$. Notice that the collections of $t$-subsets of $V(S_{C})$ which are in $\mathcal{G}_{3}$ are disjoint for distinct components $C$. We thus have that $|\mathcal{G}_{3}|\geq\alpha|\mathcal{F}_{3}|$, as required. ∎
@@ 7 Conclusion
Recall that we defined $\pi(r,t,k):=\lim_{n\to\infty}n^{-t}f^{(r)}(n;k(r-t)+t,k)$ (if the limit exists). Theorem 1.2 establishes that $\pi(r,t,k)=\frac{1}{t!}\binom{r}{t}^{-1}$ when $k$ is even and $r\geq t+(k^{3}\cdot t!)^{1/t}$. It would be interesting to prove if $\pi(r,t,k)=\frac{1}{t!}\binom{r}{t}^{-1}$ holds for every even $k$ and $r\geq t+1$, i.e. without any further restriction on $r$. We remark that this is known to be true for $k=2$ and $k=4$, as proved in [5] and [4], respectively. Concerning the case of odd $k$, Proposition 1.5 proves that $\limsup_{n\to\infty}\pi(r,t,k)\leq\frac{1}{t!}\cdot\frac{2}{2\binom{r}{t}-1}$ when $k$ is odd and $r$ is large. We wonder if this upper bound is optimal for every odd $k$, and in particular the limit $\pi(r,t,k)$ exists, given that this is the case for $k=3$, as proved by [4].

### 09 Apr 2024 v2

@@ 1 Introduction
Finally, we settle Conjecture 1.1 for $k=7$. (Our proof does not work when $r=3$ and $t=2$, but this case has been resolved for all $k$ in [2].)
@@
More about the case of $k$ being odd can be found in the concluding remarks.
Remark 1.5.
Shortly after this paper appeared on arXiv, Glock, Kim, Lichev, Pikhurko and Sun [5] determined the value of $\pi(r,t,k)$ for $t=2$, $k\in\{5,6,7\}$ and all $r\geq 3$: they showed that for $k\in\{5,7\}$, it holds that $\pi(r,2,k)=1/(r^{2}-r-1)$ (observe this is the same value as for $k=3$), while for $k=6$ it holds that $\pi(3,2,6)=61/330$ and $\pi(r,2,6)=1/(r^{2}-r)$ for $r\geq 4$ (observe this extends Theorem 1.2 when $t=2$ and $k=6$).
@@
Section 2 introduces the relevant notation and collects some preliminary results, including our key proposition (Proposition 2.4) needed for the proofs of Theorems 1.3 and 1.4, which are proved in Section 4 and Section 5, respectively. Section 3 provides the proof of Theorem 1.2 and finally Section 6 contains some concluding remarks.
@@ 2.2 Density argument
The approach of Delcourt and Postle [2], while proving Conjecture 1.1 for $r=3$, $t=2$ and any $k\geq 2$, relies on the following reduction: they show that in any sufficiently dense $k$-free $3$-graph, it is possible to find a subgraph with almost the same density which is additionally $\ell^{-}$-free for every $\ell\in[2,k-1]$, i.e. having $\ell^{-}$-configurations is ‘inefficient’ for the extremal $k$-free graph. Here we provide another density-type argument, which we will use in the proof of Theorems 1.3 and 1.4. Before stating the result, we introduce some notation. Given an $r$-graph $\mathcal{F}$, define $J(\mathcal{F})$ to be the $t$-graph with $V(\mathcal{F})$ as vertex set and where a $t$-subset $T\subseteq V(\mathcal{F})$ is an edge of $J(\mathcal{F})$ if and only if there exists an $\ell$-configuration for some $\ell\in\left[\lfloor k/2\rfloor\right]$ whose vertex set contains $T$. Observe that, since every edge is a $1$-configuration, $J(\mathcal{F})$ contains the $t$-shadow of $\mathcal{F}$. Therefore $J(\mathcal{F})$ is a supporting $t$-graph of $\mathcal{F}$ (recall Definition 2.1). Moreover, its non-edge girth is greater than $\lfloor k/2\rfloor$.
@@ 2.3 A useful lemma
We show that we can get a subhypergraph of $\mathcal{F}$ which satisfies (P1), (P2) and (P3) by removing $O(n^{t-1})$ edges, which in turn will imply (P4) as well. Observe that, since $\mathcal{F}$ is $k$-free, $\mathcal{F}$ is also $k^{-}$-free. Moreover we show that $\mathcal{F}$ can be made $(k-1)^{-}$-free by removing $O(n^{t-1})$ edges. Let $\mathcal{S}$ be a maximal collection of pairwise edge-disjoint $(k-1)^{-}$-configurations of $\mathcal{F}$. If $|\mathcal{S}|>\binom{n}{t-1}$, then there exists a set $T\subseteq V(\mathcal{F})$ of size $t-1$ which is contained in the $(t-1)$-shadow of two $(k-1)^{-}$-configurations $S_{1}$ and $S_{2}$ in $\mathcal{S}$. Let $e\in S_{2}$ be an edge such that $T\subseteq e$. Then $S_{1}\cup\{e\}$ is a $k$-configuration in $\mathcal{F}$, being a collection of $k$ edges spanning at most $[(k-1)(r-t)+t-1]+(r-|T|)=k(r-t)+t$ vertices, a contradiction to $\mathcal{F}$ being $k$-free. Therefore $|\mathcal{S}|\leq\binom{n}{t-1}$ and, by removing from $\mathcal{F}$ all edges of each $S\in\mathcal{S}$, we obtain a subhypergraph $\mathcal{F}_{0}\subseteq\mathcal{F}$ which is $(k-1)^{-}$-free and satisfies $|\mathcal{F}_{0}|\geq|\mathcal{F}|-O(n^{t-1})$.
@@
In this section, we prove Theorem 1.2, which asserts that $\lim_{n\to\infty}n^{-t}f^{(r)}(n;k(r-t)+t,k)=\frac{1}{t!}\binom{r}{t}^{-1}$, for $k$ even and $r$ sufficiently large in terms of $t$ and $k$. We do that by showing that $\liminf_{n\to\infty}n^{-t}f^{(r)}(n;k(r-t)+t,k)\geq\frac{1}{t!\,\binom{r}{t}}$, which follows directly from Corollary 2.3, and that $\limsup_{n\to\infty}n^{-t}f^{(r)}(n;k(r-t)+t,k)\leq\frac{1}{t!\,\binom{r}{t}}$, which follows from Lemma 2.6 and the inequality and claims provided below.
@@
We are now ready to prove Theorem 1.2.
@@
In this section we prove Theorem 1.3, asserting that the limit $\lim_{n\to\infty}n^{-t}f^{(r)}(n;5(r-t)+5)$ exists, for $2\leq t<r$. Our proof uses our density argument (Proposition 2.4). We recall that $J(\mathcal{F})$ is the $t$-graph on $V(\mathcal{F})$ whose edges are $t$-subsets of $\ell$-configurations in $\mathcal{F}$ with $\ell\leq\lfloor k/2\rfloor$ (this is defined above Proposition 2.4).
@@
Observe that, by definition of $J(S)$, we have that $T\in J(S)$ if and only there exists an edge $e\in S$ with $T\subseteq e$ or there exists a $2$-configuration $S^{\prime}$ in $S$ whose vertex set contains $T$. Since in $\mathcal{G}$ any $2$-configuration and any $3^{-}$-configuration are edge-disjoint, we can rule out the second option. Moreover, a set of size $t$ cannot be in more than one edge of $S$ as, otherwise, $S$ would contain a $2$-configuration, which cannot happen for the same reason. Since $S$ has three edges, we conclude that $|J(S)|=3\binom{r}{t}$. Let $T\in J(S)$. Then clearly $T\in J(\mathcal{G})$, and we aim to show that $T\not\in J(\mathcal{G}^{\prime})$. For that, note that $T\in J(\mathcal{G}^{\prime})$ if and only if there exists an edge $e\in\mathcal{G}^{\prime}$ with $T\subseteq e$ or there exists a $2$-configuration $S^{\prime}$ of $\mathcal{G}^{\prime}$ whose vertex set contains $T$. The first option cannot happen as, otherwise, $S\cup\{e\}$ would be a $4^{-}$-configuration of $\mathcal{G}$, being a collection of four edges spanning at most $(3r-2t-1)+r-t=4(r-t)+t-1$ vertices, a contradiction to $\mathcal{F}_{1}$ being $4^{-}$-free. Similarly, we can rule out the second option as, otherwise, using that $S$ and $S^{\prime}$ are edge-disjoint, $S\cup S^{\prime}$ would be a $5^{-}$-configuration of $\mathcal{G}$, being a collection of five edges spanning at most $(3r-2t-1)+(2r-t)-t=5(r-t)+t-1$ vertices, a contradiction to $\mathcal{F}_{1}$ being $5^{-}$-free.
@@
This section is concerned with the proof of Theorem 1.4, which asserts that the limit $\lim_{n\to\infty}n^{-t}f^{(r)}(n;7(r-t)+t,7)$ exists for $r>t\geq 2$ and $(r,t)\neq(3,2)$. We use our density argument (Proposition 2.4), together with the following two inequalities.
@@
We are now ready to prove Theorem 1.4.
@@ 6 Conclusion
Recall that we defined $\pi(r,t,k):=\lim_{n\to\infty}n^{-t}f^{(r)}(n;k(r-t)+t,k)$ (if the limit exists). Theorem 1.2 establishes that $\pi(r,t,k)=\frac{1}{t!}\binom{r}{t}^{-1}$ when $k$ is even and $r\geq t+(k^{3}\cdot t!)^{1/t}$. It would be interesting to determine, for fixed even $k$, what is the smallest $r$ such that $\pi(r,t,k)=\frac{1}{t!}\binom{r}{t}^{-1}$ for all $12\leq t\leq r-$ 1. We remark that the smallest such $r$ is $2$ for $k=2$ and $4$ for $k=4$, as proved in [6] and [4], respectively.
For general odd $k$ we were not able to prove that the limit $\pi(r,t,k)$ exists, even when $r$ is large. Nevertheless, arguments reminiscent of Theorem 1.2 yield that if $k$ is odd, $t\geq 2$ and $r$ is sufficiently large with respect to $k$ and $t$, then $\limsup_{n\to\infty}n^{-t}f^{(r)}(n;k(r-t)+t,k)\leq\frac{1}{t!}\cdot\frac{2}{2\binom{r}{t}-1}$. We briefly sketch the proof idea.
A $t$-tight component is a collection of edges that can be ordered as $\{e_{1},\ldots,e_{m}\}$ so that $e_{i+1}$ shares at least $t$ vertices with one of $e_{1},\ldots,e_{i}$, for each $i\in[m-1]$. Given any $k$-free $n$-vertex $r$-graph $\mathcal{F}$, we let, for $i\in[2]$, $\mathcal{F}_{i}$ be the set of the edges of $\mathcal{F}$ which belong to components of size $i$, and $\mathcal{F}_{3}:=\mathcal{F}\setminus(\mathcal{F}_{1}\cup\mathcal{F}_{2})$. For $i\in[2]$, define $\mathcal{G}_{i}$ to be the $t$-shadow of $\mathcal{F}_{i}$, and define $\mathcal{G}_{3}$ to be the collection of $t$-sets $T$ such that $T\not\in\mathcal{G}_{1}\cup\mathcal{G}_{2}$ and there is a unique component in $\mathcal{F}_{3}$ that contains a $2$-configuration whose vertex set contains $T$. With $\alpha:=\frac{1}{2}\cdot\left(2\binom{r}{t}-1\right)$, it is not hard to see that $|\mathcal{G}_{i}|\geq\alpha|\mathcal{F}_{i}|$ for $i\in[3]$. Therefore, since the sets $\mathcal{F}_{1},\mathcal{F}_{2},\mathcal{F}_{3}$ partition $\mathcal{F}$, it follows that $|\mathcal{F}|=|\mathcal{F}_{1}|+|\mathcal{F}_{2}|+|\mathcal{F}_{3}|\leq\frac{1}{\alpha}\left(|\mathcal{G}_{1}|+|\mathcal{G}_{2}|+|\mathcal{G}_{3}|\right)\leq\frac{1}{\alpha}\binom{n}{t}$, which implies the desired result.
We suspect this upper bound might be optimal, as this is the case for $k=3$ (see [4]).
Remark 6.1.
The recent work [5] mentioned in Remark 1.5 shows that, for $k=6$, the smallest $r$ such that $\pi(r,t,6)=\frac{1}{t!}\binom{r}{t}^{-1}$ for all $12\leq t\leq r-$ 1 is $r=4$. Moreover, it shows that the upper bound discussed above is optimal for $k\in\{5,7\}$.